A car can be stopped from initial velocity 84 km/h to rest in 55 meters. Assuming constant acceleration, find the stopping time.
is it 4.72 seconds?

Answer by Alan3354(69443)

A car can be stopped from initial velocity 84 km/h to rest in 55 meters. Assuming constant acceleration, find the stopping time.
is it 4.72 seconds?
---------------
Its average speed during the acceleration = 42 km/hr
42 km/hr = 42000/3600 m/sed = 35/3 m/sec
--------------
55 m/(35/3) m/sec =~ 4.714 seconds